OpenStudy (truwhovian11):

medal fan and testimonial for the first person to help me. What is the distance between -13 and 13 on a number line? A) -26 B) -13 C) 0 D) 26

OpenStudy (fortytherapper):

I don't know if this sounds to physics-like, but they're asking for the magnitude basically

OpenStudy (ineedhelplz):

I would go for D....

OpenStudy (deercult):

First thing, you can eliminate the first two answers because when it comes to distance, there are no negatives! You're basically counting the space between the numbers on the number line.

OpenStudy (truwhovian11):

so its D

OpenStudy (deercult):

Yes!

OpenStudy (ineedhelplz):

Basically \[13+(-13)=0\]\[0+13=13\]\[13+13=26\]

OpenStudy (truwhovian11):

thanks medal for @deercult and fan for all who answerd